A manufacturer purchases two machines A and B. The probability that A will last 5 years is 4/5
and the probability that B will last 5 years is 3/4.
Find the probability that:
1)both machines will last 5 years
2)only machine a will last 5 years
3) at least one of the machine will last 5 years

Answer by Boreal(15235)

a. is the product if they are independent. That is (4/5)(3/4))=3/5
-
b. That is (4/5)(1/4)=1/5, the (1/4) is probability B does not last 5 years.
-
c. At least one will last 5 years is 1- prob (both don't)
=1-(1/5)(1/4)
=19/20